x= (kyz)/p^2 what % change occurs in x, if y increases by 20%, z decreases by 20% and p is doubled
Answer by ptaylor(2198)


x= (kyz)/p^2
y increases by 20%=y+0.2y
z decreases by 20%=z-0.2z
p doubles=2p
x1={k(y+0.2y)(z-0.2z)}/(2p)^2 expand (y+0.2y)(z-0.2z) using FOIL
x1=k(yz-0.2yz+0.2yz-0.04yz)/(2p)^2 simplify
x1=(kyz-0.04kyz)/4p^2;
x1=0.96kyz/4p2;
x1=(0.96/4)(kyz/p^2)=(0.96/4)(x)
x1=0.24x
So x decreases by x-0.24x=76%
Hope this helps---ptaylor